A particle moves along the curve y=1/3x^3+2 where x and y are in meters.When x=.8, the x component of the velocity is 10m/s what is the total velocity
1
Expert's answer
2019-04-16T09:49:49-0400
"\frac{dy}{dx}=x^2"

"\frac{dy}{dt} =\frac{dy}{dx} \frac{dx}{dt} =x^2 \frac{dx}{dt}"



"\frac{dy}{dt} (0.8)=0.8^2(10)=6.4 \frac{m}{s}"

Total velocity:


"(10, 6.4) \frac{m}{s}"